The results of a Nationwide aptitude test in mathematics are normally distributed with m=80 and s=15 what percentile rank of a score of 90

1
Expert's answer
2022-03-23T17:01:03-0400

"P(X<90)=P(Z<\frac{90-80}{15})=P(Z<0.67)=0.7486."

Percentile rank: 74.